OpenStudy (anonymous):

What is the slope-intercept form of the equation 4x − y = -1?

OpenStudy (anonymous):

Take the equation and put it into the form y=mx+b by subtracting 4x nd moving it to the otherside and then divide the otherside by -y.

OpenStudy (anonymous):

y = 4x − 1 B. y = x + 1 C. y = 4x + 1 D. y = -4x + 1

OpenStudy (anonymous):

C

OpenStudy (anonymous):

C